Mr. Simon invests a total of $7003 in two savings accounts. One account yields 7.5% simple interest and the other 8.5% simple interest. He would like to find the amount placed in each account if a total of $582.33 in interest is received after one year. Write a system of linear equations to express this problem.
Answer by mananth(16949)

TOTAL INVESTMENT GIVEN. FIND INDIVIDUAL AMOUNT
7.50% per annum ------------- x
8.50% per annum ------------ y
total investment $7003
Interest----- $582.33

..
x+y= 7003 ------------------------1
7.50%x +8.50%y =$582.33
Multiply by 100
7.5x+8.5y=$58,233.00--------2

Multiply (1) by -7.5
we get

-7.5x-7.5y=-52522.5

Add this to (2)

0+1y=$5,710.50



y=$5,710.50 investment at 8.50%
Balance $1,292.50investment at 7.50%

m.ananth@hotmail.ca